OpenStudy (anonymous):

Help Please 3: Is the number 29 prime or composite? A. prime B. composite

OpenStudy (anonymous):

@Jaynator495 Please Help :3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (amistre64):

try dividing it by primes less than its closest sqrt if no primes divide into it, then itself is prime

OpenStudy (anonymous):

I think that it's Prime..

OpenStudy (anonymous):

Idk.

OpenStudy (amistre64):

29 is a small number, so test the primes less then say 7 to see if they divide into it

OpenStudy (anonymous):

\[ 29\div?\]

OpenStudy (amistre64):

2,3,5,7 .... its not even, or end in 5 or 0 .. so 3 or 7?

OpenStudy (anonymous):

Uh, i'm very sorry but i'm confused

OpenStudy (amistre64):

is it divisible by 3 or 7? it is not even so it is not divisible by 2 it does not end in 5 or 0, so it is not divisible by 5 the only primes left to test that could divide into it are 3 and 7 at best test them out

OpenStudy (anonymous):

Ok.

OpenStudy (skullpatrol):

A prime number, or prime, is an integer greater than one that has no integral factors other than itself and one.

OpenStudy (anonymous):

um, so it's composite?

OpenStudy (amistre64):

\[\sqrt{25}<\sqrt{29}<\sqrt{36}\] \[5<\sqrt{29}<6\] since sqrt(29) is no more than 6, then the primes less then 6 are needed to test 2,3,5, and 7 is just a better safe than not is 29 divisible by 2,3,5, or 7?

OpenStudy (amistre64):

is 29, divided by 2,3,5, or 7 result in an integer?

OpenStudy (anonymous):

Yeah...

OpenStudy (amistre64):

what is it divisible by?

OpenStudy (anonymous):

i don't know! I'm having trouble with this subject.

OpenStudy (skullpatrol):

A composite number has other integral factors than itself and one. It can be composed of other factors :-)

OpenStudy (amistre64):

well, i cant divide it for you ... there are 3 or 4 possible tests to run if you must.

OpenStudy (skullpatrol):

This is another way of saying that prime numbers can only be evenly divisible by 1 and them self.

OpenStudy (skullpatrol):

The answer is B, 29 is prime because no integer greater than 1 and less than 29 will divide into it evenly without a remainder.

OpenStudy (anonymous):

Yeah, i figured it out lol.

OpenStudy (skullpatrol):

Good job :-)